# VITA: Zero-Shot Value Functions via Test-Time Adaptation of Vision–Language Models

### v4

@@ D.2 Test-Time Training Hyperparameters
At inference time, we adapt only the temporal adaptation module $f_{\text{adapt}}$, using the same projection dimension ($d^{\prime}=64$) as in training. We perform a single gradient step ($t_{\mathrm{ep}}=1$) with a learning rate $\eta=0.1$. This configuration was selected from a sweep over learning rates $\{5.0,1.0,0.1,0.05,0.01,0.001\}$ and adaptation steps $\{1,5,10\}$, based on validation performance reported in Table 4. The results shows that both (i) a single update with a larger learning rate ($\eta=0.1,\,t_{\mathrm{ep}}=1$) and (ii) multiple updates with a smaller learning rate ($\eta=0.01,\,t_{\mathrm{ep}}=10$) achieve comparable validation performance under test-time training. We select the former because it requires fewer adaptation steps while achieving slightly better results.
@@ E.2 Evaluation Protocol
Following Agarwal et al. [agarwal2021deep], we report the interquartile mean (IQM) of returns across all task–seed pairs. The IQM computes the mean of the middle 50% of scores, which reduces sensitivity to outliers compared to the mean or median. For statistical robustness, we estimate 95% confidence intervals using stratified bootstrap with 10,000 resamples, stratified by task. This procedure is applied consistently across all methods in Meta-World MT10 experiments.
@@ F.1 Effect of Dissimilarity-Based Sampling on Discriminative Performance
Table 5: Disc@VOC results across scripted datasets under different training-time sampling strategies for test-time adaptation. A checkmark indicates successful discrimination between expert and suboptimal trajectories Each method is denoted as Type($w_{\text{train}}$, $k$), where: FT = full-trajectory, Rand = random sampling, Diss = dissimilarity-based sampling.
@@
We empirically validate the effectiveness of the proposed dissimilarity-based sampling by comparing the impact of different sub-trajectory sampling strategies. In particular, we evaluate how different sub-trajectory sampling strategies during training impact the ability of our model to discriminate expert from suboptimal trajectories. In Table 5, we compute Disc@VOC for 5 scripted datasets, capturing whether a given model assigns a higher average VOC to expert trajectories (tk_pnp) than to each evaluation dataset. We compare three strategies: (1) Full-trajectory sampling applies adaptation over all overlapping sub-trajectories of length $w_{\text{tr}}$ using stride 1, following prior work [sun2024rnn]; (2) Random sampling uniformly samples $k$ sub-trajectories of length $w_{\text{tr}}$ per video; and (3) our proposed Dissimilarity-based sampling constructs a candidate set using stride $\lfloor w_{\text{tr}}/2\rfloor$, and selects $k$ sub-trajectories that maximize pairwise dissimilarity in the representation space.
@@ F.2 Effect of Temporal Memory on Value Function Estimation
Shift Dataset TTT-TR TTT-VIS TTT-EX VITA
@@
We evaluate four adaptation strategies: (1) TTT-TR, which performs a single gradient update computed over the full trajectory (offline). TTT-TR performs a batched update that averages the loss across all frames, thereby discarding temporal order; (2) TTT-RS, which resets the adaptation module at every step and updates using only the current visual observation (memoryless). TTT-RS prevents any temporal information from being carried forward; (3) TTT-EX, which resets the adaptation module at every step and updates using a local window of recent frames matching the window size used during meta-training ($r=8$). TTT-EX averages gradients over this window, discarding the temporal ordering within it; (4) VITA (Ours), which updates the adaptation module incrementally at every timestep (implicit memory). VITA forms an implicit memory through the sequential accumulation of temporal history in the adapter parameters.
We performed the same test-time hyperparameter sweeps for all variants, as described in Appendix D. We found that VITA, TTT-RS, and TTT-TR achieve optimal performance with a learning rate of $0.1$ and a single adaptation step, whereas TTT-EX performs best with a learning rate of $1.0$ and one adaptation step. In the implicit-memory setting, updating over a local window of frames while not resetting would cause the most recent observation to be incorporated twice—explicitly through the window and implicitly through the accumulated parameters. Therefore, implicit memory requires sequential, per-timestep updates without reset.
The results in Table 6 show that TTT-TR, TTT-VIS, and TTT-EX yield comparable performance. This indicates that performing batched updates—whether over a short explicit window (TTT-EX) or over the full trajectory (TTT-TR)—performs similarly to the memoryless baseline (TTT-RS). In contrast, VITA consistently outperforms all three variants, underscoring the importance of sequential, per-timestep updates for effective value function estimation in real-world robotic manipulation tasks. In implicit memory (VITA), updating over a local window of frames would cause the most recent observations in that window to be incorporated twice—explicitly through the window and implicitly through the accumulated parameters. In addition, updating less frequently (non-sequentially) would cause information loss because intermediate frames are omitted. Therefore, the implicit-memory test-time training paradigm requires sequential, per-timestep updates without reset.

### v5

@@ D.2 Test-Time Training Hyperparameters
At inference time, we adapt only the temporal adaptation module $f_{\text{adapt}}$, using the same projection dimension ($d^{\prime}=64$) as in training. We perform a single gradient step ($t_{\mathrm{ep}}=1$) with a learning rate $\eta=0.1$. This configuration was selected from a sweep over learning rates $\{5.0,1.0,0.1,0.05,0.01,0.001\}$ and adaptation steps $\{1,5,10\}$, based on validation performance reported in Table 4. The results show that both (i) a single update with a larger learning rate ($\eta=0.1,\,t_{\mathrm{ep}}=1$) and (ii) multiple updates with a smaller learning rate ($\eta=0.01,\,t_{\mathrm{ep}}=10$) achieve comparable validation performance under test-time training. We select the former because it requires fewer adaptation steps while achieving slightly better results.
@@ E.2 Evaluation Protocol
Following Agarwal et al. (Agarwal et al., 2021), we report the interquartile mean (IQM) of returns across all task–seed pairs. The IQM computes the mean of the middle 50% of scores, which reduces sensitivity to outliers compared to the mean or median. For statistical robustness, we estimate 95% confidence intervals using stratified bootstrap with 10,000 resamples, stratified by task. This procedure is applied consistently across all methods in Meta-World MT10 experiments.
@@ F.1 Effect of Dissimilarity-Based Sampling on Discriminative Performance
Table 5: Disc@VOC results across scripted datasets under different training-time sampling strategies for test-time adaptation. A checkmark indicates successful discrimination between expert and suboptimal trajectories. Sampling methods FT (full-trajectory), Rand (random sampling), and Diss (dissimilarity-based sampling) are parameterized by $(w_{\text{tr}},k)$.
@@
We empirically validate the effectiveness of the proposed dissimilarity-based sampling by comparing the impact of different sub-trajectory sampling strategies. In particular, we evaluate how different sub-trajectory sampling strategies during training impact the ability of our model to discriminate expert from suboptimal trajectories. In Table 5, we compute Disc@VOC for 5 scripted datasets, capturing whether a given model assigns a higher average VOC to expert trajectories (tk_pnp) than to each evaluation dataset. We compare three strategies: (1) full-trajectory sampling applies adaptation over all overlapping sub-trajectories of length $w_{\text{tr}}$ using stride $s$, following prior work (Sun et al., 2024); (2) random sampling uniformly samples $k$ sub-trajectories of length $w_{\text{tr}}$ per video; and (3) our proposed dissimilarity-based sampling constructs a candidate set of sub-trajectories of length $w_{\text{tr}}$ using stride $s$, and selects $k$ sub-trajectories that maximize pairwise dissimilarity in the representation space.
@@ F.2 Effect of Temporal Memory on Value Function Estimation
Shift Dataset TTT-TR TTT-RS TTT-EX VITA
@@
We evaluate four adaptation strategies: (1) TTT-TR, which performs a single gradient update computed over the full trajectory (offline). TTT-TR performs a batched update that averages the loss across all frames, thereby discarding temporal order; (2) TTT-RS, which resets the adaptation module at every step and updates using only the current visual observation (memoryless). TTT-RS prevents any temporal information from being carried forward; (3) TTT-EX, which resets the adaptation module at every step and updates using a local window of recent frames matching the window size used during meta-training ($w_{\text{tr}}=8$). TTT-EX averages gradients over this window, discarding the temporal ordering within it; (4) VITA (Ours), which updates the adaptation module incrementally at every timestep (implicit memory). VITA forms an implicit memory through the sequential accumulation of temporal history in the adapter parameters. We performed the same test-time hyperparameter sweeps for all variants, as described in Appendix D. We found that VITA, TTT-RS, and TTT-TR achieve optimal performance with a learning rate of $0.1$ and a single adaptation step, whereas TTT-EX performs best with a learning rate of $1.0$ and one adaptation step.
The results in Table 6 show that TTT-TR, TTT-RS, and TTT-EX yield comparable performance. This indicates that performing batched updates—whether over a short explicit window (TTT-EX) or over the full trajectory (TTT-TR)—performs similarly to the memoryless baseline (TTT-RS). In contrast, VITA consistently outperforms all three variants, underscoring the importance of sequential, per-timestep updates for effective value function estimation in real-world robotic manipulation tasks. In implicit memory (VITA), updating over a local window of frames would cause the most recent observations in that window to be incorporated twice—explicitly through the window and implicitly through the accumulated parameters. In addition, updating less frequently (non-sequentially) would cause information loss because intermediate frames are omitted. Therefore, the implicit-memory test-time training paradigm requires sequential, per-timestep updates without reset.